A turbine shaft of diameter 250 mm is running at 1600 r.p.m. in a journal bearing and supports a load of 130 kN. Calculate (a) length of bearing if the permissible bearing pressure is 1.5 N/mm2 , (b) coefficient of friction, (c) rubbing velocity, (d) amount of heat to be removed by the lubricant per minute. The bearing temperature is 58 °C and viscosity of oil at this term is 0.02 kg/m-s. The bearing clearance is 0.25 mm, k = 0.002.


1
Expert's answer
2021-08-19T01:19:02-0400

coefficient of friction,

W =130,000N

N = 1600 rpm

Ta=580C

Z = 0.02 "\frac{kg}{m-s}"

D = 250 mm = 0.25m

C =0.25 mm = 0.00025 m

k =0.002


coefficient of friction,

 

"\mu = \frac{33}{10^8}\frac{ZN}{P}\frac{d}{c}+k"


"\mu = \frac{33}{10^8}\frac{0.02\times1600}{1.25}\frac{0.25}{0.00025}+0.002 = 0.0284"

 rubbing velocity

"V = \frac{\pi DN}{60}=\frac{\pi \times 0.25\times1600}{60} =20.94 \frac{m}{sec}"


amount of heat to be removed by the lubricant per minute.

"Q_g = \mu WV =0.0284\times130,000\times20.94 =77.31 kW"